julia has a rectangular garden with a perimeter of 72m and an area of 308^2m. what is the length and width of the garden
Answer by richwmiller(17219)

2L+2w=72,
l+w=36
w=36-L
L*w=308
L*(36-L)=308
36l-l^2-308=0
-(l-22) (l-14) = 0
l=22 w=14